### v2

# Pose-dIVE: Pose-Diversified Augmentation for Person Re-Identification

Abstract
Person re-identification (Re-ID) often faces challenges due to variations in human poses and camera viewpoints, which significantly affect the appearance of individuals across images. Existing datasets frequently lack diversity and scalability in these aspects, hindering the generalization of Re-ID models to new camera systems. We propose Pose-dIVE, a novel data augmentation approach that incorporates sparse and underrepresented human pose and camera viewpoint examples into the training data, addressing the limited diversity in the original training data distribution. Our objective is to augment the training dataset to enable existing Re-ID models to learn features unbiased by human pose and camera viewpoint variations. To achieve this, we leverage the knowledge of pre-trained large-scale diffusion models. By conditioning the diffusion model on both the human pose and camera viewpoint concurrently through the SMPL model, we generate training data with diverse human poses and camera viewpoints. Experimental results demonstrate the effectiveness of our method in addressing human pose bias and enhancing the generalizability of Re-ID models compared to other data augmentation-based Re-ID approaches.
footnotetext: Equal contribution.
Introduction
Person re-identification (Re-ID) has been popularly used to retrieve (Zheng, Yang, and Tian 2017) and track (Wang 2013) individuals across images captured by multiple cameras with non-overlapping views (Chen et al. 2018; Zheng, Yang, and Hauptmann 2016). It has applications in various fields, such as robotics (Coşar and Bellotto 2020) and surveillance (Wang 2013).
However, real-world deployment of Re-ID often faces several challenges. A major obstacle arises from the differences in human pose (Cho and Yoon 2016; Sarfraz et al. 2018; Zhao et al. 2017) and camera viewpoint (Bak et al. 2014; Karanam, Li, and Radke 2015) within images of the same individual. These variations significantly affect the appearance of individuals, making the identification task difficult. To overcome this, existing models often require diverse images, but datasets for Re-ID are limited in terms of scalability and generalizability (Fu et al. 2021), since manual labeling of individuals across cameras is notoriously challenging (Liu et al. 2014; Zhao, Ouyang, and Wang 2013). Furthermore, the limited number of cameras in the datasets makes it challenging to generalize models to new camera networks (Luo, Song, and Zhang 2020; Wang 2013; Zhong et al. 2018). As a result, learning pose-invariant features for Re-ID that remain consistent across different human poses and camera viewpoints remains a challenge.
Figure 1: Pose-dIVE diversifies the viewpoint and human pose of the Re-ID dataset to help generalize and improve the performance of arbitrary Re-ID models.
Figure 2: Visualization of the effect of viewpoint and human pose augmentation. We compare visualizations of camera viewpoint and human pose distributions for the Market-1501 (Zheng et al. 2015). The left figures (i) display the camera viewpoint distribution derived from SMPL, while the right figures (ii) illustrate the pose distribution. In (i), from left to right, we show the viewpoint distributions of the training dataset, the augmented dataset, and the combination of both. Similarly, in (ii), from left to right, we present t-SNE (Van der Maaten and Hinton 2008) visualizations of the human pose distributions, showing poses from the training dataset, followed by augmented poses sourced from outside the dataset. These visualizations demonstrate that our pose augmentation successfully diversifies both viewpoint and human pose distributions.
To address these challenges, previous attempts (Dai et al. 2019; Fu et al. 2021; Gong, Huang, and Chen 2021a; Huang et al. 2018; Liu et al. 2018; Luo et al. 2019; Zheng et al. 2019) have involved data augmentation aimed at increasing the diversity of training data. Early methods employed simple image transformation techniques (Huang et al. 2018; Luo et al. 2019; Zhong et al. 2020), while more recent approaches leverage generative adversarial networks (GAN)-based models for augmentation (Liu et al. 2018; Zheng et al. 2019). However, since these methods rely on poses already existing in the Re-ID dataset to generate augmented images, they do not effectively address the core issue of pose bias in the training data distribution. Even in the case that the human pose is the same, differences in camera angles may induce significant changes in appearance (Ye et al. 2021). For example, a picture taken from a high-angle camera viewpoint may appear different from one taken from a low-angle viewpoint, potentially causing the same person to be identified as different individuals. To generalize and reduce the training data bias, it is crucial to diversify the training data’s human pose and camera viewpoint distribution.
In this paper, we define a novel pose augmentation technique which enables Re-ID models to robustly handle pose (i.e., human pose and camera viewpoint) variation, as illustrated in Figure 1. We propose a diffusion model-based framework, called Pose-dIVE, to augment the training dataset with generated images that contain human poses and camera viewpoints relatively underrepresented (i.e., rarely appearing) within the original training dataset. This effectively normalizes the distribution of the training dataset, allowing Re-ID models trained to be more generalizable and scalable.
However, a generative model naively trained under a Re-ID dataset would struggle to generate such images with infrequent poses and cameras, as such a network would likewise reflect the bias within the training data distribution. To overcome this issue, we leverage Stable Diffusion (Rombach et al. 2022), a pre-trained large-scale generative model, whose powerful generative priors enable realistic, robust generation of images with human pose and camera conditions uncommon in the training dataset. Also, to jointly provide human pose and camera viewpoint that are rarely seen in the training dataset, we use the SMPL (Loper et al. 2015) model. We condition the generative model with SMPL, enabling realistic human rendering and providing implicit information about the camera viewpoint through depth data.
Our method focuses on augmenting a training dataset in a manner that enables Re-ID models to learn features unbiased to pose variations. We present performance gains achieved by training the Re-ID model on our offline augmented dataset using our proposed architecture.
This highlights the effectiveness of our method in addressing the pose bias issue and enhancing the generalizability of Re-ID models.
Figure 3: Pose-dIVE framework. Upon observing the highly biased viewpoint and human pose distributions in training dataset, we augment the dataset by manipulating SMPL body shapes and feeding the rendered shapes into a generative model to fill in sparsely distributed poses and viewpoints. With this augmented dataset, we can train a Re-ID model that is robust to viewpoint and human pose biases.
Related Work
Person re-identification.
Person Re-ID is a crucial task in computer vision that involves identifying individuals across different camera views. Recent advancements in Re-ID (Chen et al. 2023; Fu et al. 2021; Gong, Huang, and Chen 2021b; He et al. 2021b; Kapil 2021; Li, Sun, and Li 2023; Somers, De Vleeschouwer, and Alahi 2023; Wang et al. 2019; Wieczorek, Rychalska, and Dabrowski 2021) leverage deep learning techniques to improve Re-ID performance. The most significant method is Distance Metric Learning (Hirzer et al. 2012; Koestinger et al. 2012; Liao and Li 2015; Xiong et al. 2014; Yu, Wu, and Zheng 2018; Zheng, Gong, and Xiang 2011), such as Triplet Loss (Hoffer and Ailon 2015), which utilizes deep learning models to project images into a feature embedding space where the vectors of the same identity are close together, while those of different identities are far apart.
Data augmentation in re-identification.
Various approaches (Dai et al. 2019; Luo et al. 2019; Huang et al. 2018) employed data augmentation techniques to enhance person Re-ID models. Operations such as random resizing, cropping, and horizontal flipping (Luo et al. 2019), along with adversarially occluded samples (Huang et al. 2018), increase the diversity of training data. Approaches using methods such as DropBlock (Dai et al. 2019) and background transformation (McLaughlin, Del Rincon, and Miller 2015) show further improved feature learning. Meanwhile, there are approaches (Zheng, Zheng, and Yang 2017; Liu et al. 2018; Zheng et al. 2019) using a generative model, which is Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), for data augmentation in Re-ID task. DCGANs (Zheng, Zheng, and Yang 2017) generate new samples, enriching training datasets which improves the performance of the Re-ID model. Liu et al. (2018) utilizes a data augmentation method involving pose-transferred sample augmentations. In Zheng et al. (2019), data augmentation is achieved through a joint learning framework that couples generative and discriminative modules.
Pose-conditioned diffusion models.
Pose-guided person image generation task aims to generate a person’s image with a specific pose. With the promising result of diffusion model (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020), there have been approaches (Bhunia et al. 2023; Zhang, Rao, and Agrawala 2023; Rombach et al. 2022; Han et al. 2023; Hu et al. 2023; Karras et al. 2023; Xu et al. 2023; Zhu et al. 2024) that aim to generate human images with desired structure by applying diffusion model with 2D human pose as a condition. Leveraging the extensive pre-trained knowledge of Stable Diffusion (Rombach et al. 2022), these methods demonstrate impressive generalizability across various modalities and human poses. In this work, our goal is not to develop a superior pose-conditioned diffusion model, but rather to propose a framework that utilizes the inherent generalizability of diffusion models for the task of Re-ID.
Method
Motivation and Overview
In this paper, we tackle the challenging Re-ID problem of limited camera viewpoints and human poses in training datasets. Re-ID datasets typically contain a limited number of cameras, ranging from 6 in the CUHK03 dataset (Li et al. 2014), 6 in Market-1501 (Zheng et al. 2015) to 15 in the MSMT17 dataset (Wei et al. 2018). Note that MSMT17 has the largest number of cameras among the publicly available Re-ID datasets. This small number of fixed cameras capture individuals within a restricted range of viewpoints, significantly impacting the generalization performance of Re-ID models. Furthermore, since the cameras mostly capture people with limited variations in human poses, such walking or standing still, this also affects the generalization capabilities of Re-ID models. The Figure 2 illustrates the restricted camera viewpoints and human poses captured in the training dataset.
These limitations in datasets have been a long-standing crucial problem in Re-ID models (Ye et al. 2021). Early research efforts attempted to solve this challenge of pose and camera variations through algorithmically or hand-crafted feature construction based on body structures (Zhong et al. 2018; Somers, De Vleeschouwer, and Alahi 2023). However, the restricted camera viewpoints and human poses are inherent issues in the dataset itself, and solely relying on algorithmic or hand-crafted methods has limitations in fully resolving these problems.
We aim to diversify the range of camera viewpoints and human poses by augmenting the training dataset with generative models. However, as generative models sample from the training dataset distribution, without explicit consideration for generalization, the quality of generated images with camera viewpoints and human poses that are rarely present in the training dataset distribution may be compromised. To address this, we leverage the powerful generalization capability of a foundation generative model trained on a billion-scale dataset, i.e, Stable Diffusion (Rombach et al. 2022). We carefully select and integrate components to enable camera viewpoint and human pose conditioning within the generative model for augmentation. By leveraging its generalization ability and applying viewpoint conditioning, we can effectively augment the training dataset with diverse camera viewpoints and human poses, as shown in Figure 3.
In this section, we first explain the augmentation process. Subsequently, we describe how human pose and camera viewpoint conditions are derived from the SMPL model, and how to incorporate these conditions along with identity information into the diffusion model.
Human Pose and Camera Viewpoint Condition with SMPL
To augment the data with varying human pose and camera viewpoint, we provide specific conditions to the generative model. Previous works (Ge et al. 2018; Qian et al. 2018; Tang et al. 2020) have primarily addressed pose control by providing a human pose skeleton. However, relying solely on the human skeleton has limitations due to missing information: when projected onto 2D images, the human skeleton lacks depth information, posing ambiguity for the model when inferring viewpoint information from the skeleton. For example, if the camera is above a person, the skeleton would appear compressed vertically. Without depth information, the model cannot distinguish whether the camera is positioned above the person or the person is simply short.
In this regard, in addition to the human pose skeleton, we utilize SMPL (Loper et al. 2023), a human body model used for realistic human rendering. SMPL can model intricate human shapes, including complex body articulations in 3D space. From this human model, we can easily extract 2D representations of the 3D human by rendering the model, such as depth maps which implicitly contain camera viewpoint information.
To enrich the camera viewpoint information, we incorporate depth maps. Additionally, surface normals from SMPL are used to capture detailed human surface characteristics, enhancing the precision of the generated augmentations. As a result, depth maps, surface normals, and human skeletons from SMPL, along with a reference image to control identity, are fed into the generative model as guidance.
Pose Diversified Augmentation
To mitigate the biased camera viewpoint and human pose in training dataset, we augment images with uniformly distributed camera viewpoint and human poses sourced from outside the training dataset. For camera viewpoints, we augment the images adjusting two factor: elevation and azimuth. For the elevation, we uniformly sample from $0^{\circ}$ to $30^{\circ}$, assuming that the camera is not positioned below the ground and that a person becomes indistinguishable when the camera is positioned above a certain degree. For the azimuth, we randomly sample from $0^{\circ}$ to $360^{\circ}$, as a person can be captured in any random direction. With these considerations, we found the distribution of camera viewpoint is significantly unbiased, as depicted in Figure 2. For human poses, we source them from outside the training dataset for diversification, e.g., using poses extracted from dance videos (Chan et al. 2019). As Re-ID datasets often contain similar human poses, solely training on these datasets can lead to overfitting and limited generalization capabilities. By incorporating these diverse external poses, our approach enhances the model’s ability to handle unseen poses, thereby addressing this limitation and improving overall performance. Using these diversified viewpoint and pose distributions, we render SMPL models into 2D representations to serve as conditions for the generative model.
Pose Diversified Generation with Stable Diffusion
We generate training images with diverse camera viewpoints and human poses to reduce bias in the distribution of the training data. However, when training the generative model on a human Re-ID dataset without careful consideration, it may produce degenerated results for camera viewpoints or human poses that are rarely present in the training dataset. In this work, we address this problem by leveraging the vast knowledge in pre-trained Stable Diffusion (Rombach et al. 2022).
The way we introduce human identity as a condition into the diffusion model is adapted from the framework proposed in the work by Hu et al. (2023). This approach is effective in preserving the identity of the input image while leveraging the extensive pre-trained knowledge of Stable Diffusion. The approach clones the pre-trained diffusion model into two branches of U-Nets. One branch, a reference U-Net, receives an image of a person whose identity is to be generated, while the other is a denoising U-Net that gradually removes Gaussian noise. A reference U-Net provides the identity information to the denoising U-Net through an attention mechanism. Identity information is shared with the denoising U-Net using self-attention in each block, allowing the two parallel branches to benefit from the comprehensive pre-trained knowledge of Stable Diffusion.
In addition, we concatenate the depth, surface normals, and skeleton, and feed them into the pose guider network, which consists of several convolutional layers. The encoded condition is then added to the projected input of the denoising diffusion model. In this manner, the model can generate an image with the identity given by the reference image, while being able to control its viewpoint and human pose. This enables the reduction of bias in the training dataset when a Re-ID model is trained using this dataset. For a detailed description of the architecture, please refer to the supplementary materials.
Methods MSMT17 Market1501 CUHK03 (D) CUHK03 (L)
mAP $\uparrow$ R1 $\uparrow$ mAP $\uparrow$ R1 $\uparrow$ mAP $\uparrow$ R1 $\uparrow$ mAP $\uparrow$ R1 $\uparrow$
TransReID (He et al. 2021a) 69.4 86.2 89.5 95.2 - - - -
AAFormer (Zhu et al. 2023) 65.6 84.4 88.0 95.4 77.2 78.1 79.0 80.3
AGW (Ye et al. 2021) 49.3 68.3 87.8 95.1 - - 62.0 63.6
FlipReID (Ni and Rahtu 2021) 68.0 85.6 89.6 95.5 - - - -
CAL (Rao et al. 2021) 64.0 84.2 89.5 95.5 - - - -
PFD (Wang et al. 2022) 64.4 83.8 89.7 95.5 - - - -
SAN (Jin et al. 2020) 55.7 79.2 88.0 96.1 74.6 79.4 76.4 80.1
LDS (Zang et al. 2021) 67.2 86.5 90.4 95.8 - - - -
MPN (Ding et al. 2020) 62.7 83.5 90.1 96.4 79.1 83.4 81.1 85.0
MSINet (Gu et al. 2023) 59.6 81.0 89.6 95.3 - - - -
SCSN (Chen et al. 2020) 58.5 83.8 88.5 95.7 81.0 84.7 84.0 86.8
Baseline (CLIP-ReID) 68.0 85.8 89.6 95.5 93.7 95.5 95.5 96.6
+ Pose-dIVE 71.0 87.5 90.3 95.6 95.5 97.4 97.2 97.8
Baseline (SOLIDER) 67.4 85.9 91.6 96.1 95.6 96.7 97.4 98.5
+ Pose-dIVE 68.3 85.9 92.3 96.6 96.2 97.2 97.6 98.5
Table 1: Quantitative comparison on standard Re-ID benchmarks. Since our generative augmentation can be applied to any Re-ID model, we trained two recent state-of-the-art baselines (Li, Sun, and Li 2023; Chen et al. 2023) with Pose-dIVE.
Experiments
Implementation Details
We use CLIP-reID (Li, Sun, and Li 2023) and SOLIDER (Chen et al. 2023) as the baseline for validate the effectiveness of the Pose-dIVE augmented dataset. The entire training process is divided into three parts: training the generative model, generating images for augmentation, and training the baseline Re-ID models using the augmented dataset.
Step 1: Training of generative model.
The training of our generative model involves two stages. First, it learns a general human representation using a fashion video dataset (Zablotskaia et al. 2019), followed by fine-tuning with person Re-ID datasets. Throughout both stages, the weights of the autoencoders (Yu et al. 2021) and the CLIP (Radford et al. 2021) image encoder are kept frozen, focusing on the learning of reference U-Net, denoising U-Net, and pose guider. Initially, the reference and denoising U-Nets are initialized from the Stable Diffusion (Rombach et al. 2022) model and fine-tuned with the fashion video dataset. We employ Mean Squared Error (MSE) loss, optimized using Adam (Kingma and Ba 2014) with a learning rate of 1e-5 and weight decay of 0.01. The first stage takes approximately 15 hours on a single NVIDIA RTX A6000 GPU with a batch size of 2. In the second stage, the model is further fine-tuned using Re-ID datasets consisting of pedestrian images from CCTV cameras, with the images resized to 192$\times$384 and the batch size increased to 4.
Step 2: Augmentation with generative model.
Our approach leverages data augmentation via a generative model to enrich the training data for the Re-ID model. We begin by rendering SMPLs extracted from the Everybody Dance Now dataset (Chan et al. 2019), an external dataset not used during Re-ID model training. From these SMPLs, we generate skeleton, depth, and normal maps using various camera viewpoints and a wide range of human poses, creating a “target condition gallery.” For each person identity (PID), we randomly select one instance and pair it with a randomly chosen target condition from this gallery. This pair is then fed into the generative model to produce a new image. We repeat this process iteratively, ensuring the same number of iterations for each PID. These generated images are then combined with the original training datasets for training.
Step 3: Training baseline Re-ID models.
We trained two baseline Re-ID models, CLIP-reID and SOLIDER, using both the original and augmented Re-ID datasets. For training the baselines, we followed the procedure outlined in their respective papers.
Evaluation Protocol
We evaluate the performance of the models using the four publicly available person Re-ID datasets, i.e., MSMT17 (Wei et al. 2018), Market-1501 (Zheng et al. 2015), CUHK03 (L), and CUHK03 (D) (Li et al. 2014). L and D stands for Labeled and Detected, respectively. For evaluation metrics, we use two standard Re-ID metrics: cumulative matching characteristics at Rank-1 (R1) and mean average precision (mAP).
Figure 4: Qualitative comparison. We compare our generated output with DPTN (Zhang et al. 2022), showing that Pose-dIVE can generate more realistic images while better preserving identity and accurately following the target pose.
Human Pose Viewpoint MSMT17 Market-1501 CUHK03 (D) CUHK03 (L)
Augmentation Augmentation mAP R1 mAP R1 mAP R1 mAP R1
(I) 68.0 85.8 89.6 95.5 93.7 95.5 95.5 96.6
(II) 70.9 87.0 90.1 95.3 93.8 95.3 95.8 96.8
(III) 70.9 87.3 90.2 95.4 94.6 96.6 96.4 97.4
(IV) 71.0 87.5 90.3 95.6 95.5 97.4 97.2 97.8
Table 2: Quantitative validation of Pose-dIVE augmentation strategies.
Figure 5: Qualitative results. Example images from the augmented MSMT17 and Market-1501 dataset demonstrate how the generated images preserve original identities while maintaining realism and consistency with the Re-ID dataset.
Figure 6: Visualization of the split data. To validate the generalization power of our framework, we split the MSMT17 dataset into train/test sets using two distinct approaches: 1) splitting based on viewpoint, and 2) splitting based on human pose. The visualization clearly illustrates the separation between the train and test distributions.
Training Dataset # of Images PIDs Market1501
mAP $\uparrow$ R1 $\uparrow$
(I) Baseline Dataset 11,883 619 74.7 88.9
(II) Real Images Augmented 30,453 619 77.8 92.8
(III) MARS 495,857 619 72.4 82.3
(IV) Pose-dIVE Augmented 30,453 619 80.2 92.9
Table 3: Comparison with real-images augmentation. Note that we used the ResNet50 model for efficiency.
Quantitative Comparisons
The quantitative results, presented in Table 1, demonstrate the effectiveness of our approach. To emphasize its applicability to arbitrary Re-ID models, we conducted experiments on two state-of-the-art models (Li, Sun, and Li 2023; Chen et al. 2023). Both models exhibited significant performance boosts when trained on dataset augmented with our approach. This validates the broad applicability of our augmentation method to a variety of Re-ID models. Furthermore, we observed consistent performance improvements across various datasets, further validating our augmentation framework and highlighting the efficacy of addressing sparsely distributed human poses and viewpoints.
Qualitative Results
Qualitative comparisons.
In Figure 4, we present a qualitative comparison to a recent GAN based approach (Zhang et al. 2022). In contrast to GAN-based methods, our approach effectively generates poses sourced from outside the dataset. GAN-based methods struggle to generalize to diverse poses, often resulting in blurry outputs and a limited ability to maintain the identity of both the reference and the target pose, particularly when dealing with complex features such as accessories, specific clothing details, or bags. This underscores the advantage of utilizing a pre-trained diffusion model, which possesses significant general knowledge about the world.
Visualization of generated data.
In Figure 5, we present visualization results on two datasets, MSMT17 (Wei et al. 2018) and Market-1501 (Zheng et al. 2015). Given a reference image, our method faithfully preserves its identity while being able to generate diverse poses with high fidelity.
Method MSMT17-S (Cam) MSMT17-S (Pose)
mAP $\uparrow$ R1 $\uparrow$ mAP $\uparrow$ R1 $\uparrow$
Baseline (SOLIDER) 46.6 66.7 60.0 76.0
+ Pose-dIVE 52.9 71.1 63.7 78.4
Table 4: Generalization performance on custom split. Cam denotes dataset split with camera viewpoint, while Pose denotes dataset split with human pose. Pose-dIVE demonstrates generalization even on extreme camera viewpoints and human poses not present in the training set.
Method Market-1501
mAP $\uparrow$ R1 $\uparrow$
Baseline (CBN) 77.3 91.3
+ Pose-dIVE 82.3 93.4
Baseline (BPBreID) 89.4 95.7
+ Pose-dIVE 89.9 95.8
Table 5: Pose-dIVE with generalization-focused methods. We use Market1501 as the benchmark.
Ablation Study and Analysis
Ablation on the Pose-dIVE augmentation strategy.
In Table 2, we conduct an ablation study using CLIP-ReID (Li, Sun, and Li 2023) to verify the effectiveness of our viewpoint and human pose augmentation. (I) serves as the baseline, representing a Re-ID model trained on the original dataset without our augmentation. For (II) and (III), we augment viewpoints and human poses, respectively. Both augmentations demonstrate significant performance gains, validating the effectiveness of targeting the biased distributions of human pose and viewpoint for augmentation. (IV) demonstrates the full augmentation strategy of our model. The performance gains observed in (IV) compared to both (II) and (III) confirm that both types of augmentation are not only beneficial individually, but also exhibit a complementary effect when combined, leading to further improvements in performance.
Comparison with large real image augmented datasets.
In this analysis, we investigated whether the performance gain from augmentation originates from an increased dataset size or from increased pose diversity. To validate this, we compared our augmentation method with real-world data collection as shown in Table 3. MARS (Zheng et al. 2016) is a extended version of Market-1501, which sampling more images from the same videos used in Market-1501.
The baseline training dataset, (I), is a subset of Market-1501, filtering out identities not present in MARS. (II) is the subset of the MARS dataset, matching the number of training images with ours. (III) represents the complete MARS training dataset, which has a vast collection of images. (IV) is the dataset augmented with our approach.
In (II), when the same number of images is added to the training set, our method significantly outperforms the real images sampled without considering pose and viewpoint. In (III), even though the real-image augmented dataset (MARS) is approximately 16 times larger than our augmented dataset, the Re-ID model trained on the real-image augmented dataset exhibits lower performance in both mAP and Rank-1 (R1) metrics compared to the model trained on our augmented dataset. This emphasizes that the performance of Re-ID models is not solely determined by the size of the dataset but rather by the diversity and generalization capability of the images within the dataset.
Analysis on generalization effectiveness.
To rigorously assess the generalization capabilities of our framework, we devised an experiment involving an extreme dataset split of the MSMT17 dataset, intentionally amplifying the bias of human pose and viewpoint in the training dataset. We propose two distinct split variants: MSMT17-S (Cam) and MSMT17-S (Pose), where Cam divides the dataset into train and test sets with non-overlapping camera viewpoints, while Pose divides it based on human pose. These datasets create clear experimental conditions by intentionally partitioning the dataset according to viewpoints and human poses, as visualized in Figure 6. This figure provides a clear visual representation of how the data was split and the lack of overlap between the training and test sets for both viewpoints and human poses. For each split, we fine-tune the diffusion model on the training split and generate the dataset. Then, we use the generated dataset to train the baseline model. We select SOLIDER as the baseline model for this experiment.
As shown in Table 4, the results clearly demonstrate that our approach significantly improves generalization. The performance enhancements on these filtered datasets were substantially higher, indicating the effectiveness of our method in diversifying the viewpoints and human poses of the original dataset, thereby improving the model’s ability to generalize to new, unseen pose, even in the extreme setup. For further details on the dataset split process, please refer to the supplementary material.
Analysis on the synergy of generalization-focused Re-ID models with Pose-dIVE.
In Table 5, we explore the synergy between our approach and Re-ID models specifically designed to enhance generalization, particularly those addressing variations in camera and human poses.
We evaluate the impact of our approach on two representative models: CBN (Zhong et al. 2018), which addresses camera bias, and BPBreID (Somers, De Vleeschouwer, and Alahi 2023), designed to handle human pose variations. We select these models due to their public availability and proven effectiveness in their respective domains. While both models are already effective, they show marked improvements when trained on the dataset augmented by our approach. These results highlight that even with algorithmic advancements in pose generalization, there remains significant potential for improvement through dataset enhancements.
Method Market-1501
mAP $\uparrow$ R1 $\uparrow$
w/o Pre-trained Diffusion 82.7 93.2
w/ Pre-trained Diffusion 90.3 95.6
Table 6: Ablation on using pre-trained diffusion model. We train the CLIP-reID model (Li, Sun, and Li 2023) using the generated dataset.
Ablation on pre-trained diffusion model.
In Table 6, we conduct an ablation study to evaluate the impact of utilizing the pre-trained Stable Diffusion model (Rombach et al. 2022) for generating diverse camera viewpoints and human poses in our augmented dataset. The objective is to determine whether leveraging the model’s inherent generalization ability could effectively enhance Re-ID model performance. Employing the pre-trained diffusion model results in a notable 7.6% mAP improvement, underscoring the significance of foundation models in the context of dataset augmentation.
Conclusion
In this paper, we proposed Pose-dIVE, a novel data augmentation approach that leverages pre-trained diffusion models to diversify pose and viewpoint distributions in person re-identification training datasets. The key to Pose-dIVE’s success lies in diversifying human pose and viewpoint by integrating pose, viewpoint, and identity conditions into large-scale pre-trained diffusion models, effectively leveraging the vast knowledge embedded in these models for generating high-quality augmented data.
Comprehensive experiments demonstrated the effectiveness of our approach, with Pose-dIVE achieving significant performance improvements compared to baselines. Ablation studies further verified the importance of both diverse pose augmentation and camera viewpoint augmentation.
Appendix
Detailed Explanation of MSMT17-S
In the domain of person re-identification (Re-ID), addressing the biases within datasets are crucial for improving model generalization. Such biases can lead to models that perform well under familiar conditions but struggle when faced with new scenarios. Among the various biases present, we focused on addressing the camera viewpoint bias and human pose bias inherent in the dataset through dataset augmentation. To rigorously evaluate and challenge the generalization capabilities of our Pose-dIVE method, we create two specialized versions of the MSMT17 dataset: MSMT17-S (Cam) and MSMT17-S (Pose).
MSMT17-S (Cam) is specifically designed to amplify the bias related to camera viewpoints. In this version of the MSMT17 dataset, the training set consists of images captured from a specific set of camera angles, while the test set includes entirely different angles that do not overlap with those in the training data. This deliberate split maximizes the viewpoint bias in the training data, creating a scenario where the model must generalize to new, unseen camera angles during testing. MSMT17-S (Cam) serves as a challenging benchmark to evaluate whether the Pose-dIVE augmentation method can effectively help models overcome viewpoint biases, which are prevalent in real-world Re-ID scenarios where camera placements vary widely.
MSMT17-S (Pose), on the other hand, is designed to emphasize the bias related to human poses. In this dataset, the training data is composed of individuals captured in a specific range of poses, while the test set features entirely different poses that are not present in the training data. By maximizing pose bias in this way, MSMT17-S (Pose) creates a challenging environment where a Re-ID model must generalize to new poses that it has never seen during training. This setup is particularly relevant for testing how well Pose-dIVE can assist a model in adapting to a wide variety of human postures, which is crucial for accurate person identification across different activities and positions.
These datasets specifically highlight the challenges posed by biases in camera viewpoints and human poses, respectively, allowing us to assess the effectiveness of the Pose-dIVE augmentation strategy in mitigating these biases.
Conditional Diffusion Models for Pose-Diversified Augmentation
Figure 7: Overall architecture of generative model in Pose-dIVE. Given the viewpoint and pose distributions, we first render the body shape sampled from the distribution using SMPL, generating the corresponding skeleton, depth map, and normal maps. These conditions, along with a reference image for identity preservation, are then fed into generative module, which consists of two branches: the reference U-Net processes the identity information from the reference image, while the denoising U-Net generates a person with the same identity, given the input conditions. The denoising U-Net generates images by iterating through the denoising process.
In this section, we delineate the architecture of the conditional diffusion model we employed. This model is designed to generate images with diverse camera viewpoints and human poses, conditioned on the relevant input data. However, naively training the generative model on a human Re-ID dataset without careful consideration may produce degenerated results for camera viewpoints or human poses that are rarely present in the training dataset.
We address this problem by leveraging the vast knowledge in pre-trained Stable Diffusion (Rombach et al. 2022). We first provide a preliminary explanation of the Stable Diffusion model (Rombach et al. 2022), followed by thr method for augmenting the images in the training dataset while controlling their human pose, camera viewpoint, and identity. The overall architecture can be seen in Figure 7.
Preliminary: Stable Diffusion model.
Diffusion model (Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Song, Meng, and Ermon 2020) is a generative model that samples images from the learned data distribution $p(x)$ through iterative denoising process from Gaussian noise. Our method builds upon Stable Diffusion (SD) (Rombach et al. 2022). SD performs a denoising process in a latent space of Autoencoder (Yu et al. 2021), reducing the computational cost compared to denoising in the pixel space (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020). Specifically, the encoder in SD maps a given image x into a latent representation z, denoted as $\textbf{z}=\mathcal{E}(\textbf{x})$.
In training time, SD learns a denoising U-Net (Ronneberger, Fischer, and Brox 2015) $\epsilon_{\theta}$ that predicts normally distributed noise $\epsilon$ given a noised latent $\textbf{z}_{t}$, which is a noisy latent of z with a Gaussian noise at noise level $t$. This U-Net function can be trained with a following objective:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}_{t},c,\epsilon,t}\left(\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},c,t\right)\|^{2}_{2}\right),$ (1)
where $c$ denotes conditional information for generation. The condition, which a text prompt encoded using the CLIP text encoder (Radford et al. 2021), enables controllability over the image generation process. The denoising U-Net is composed of three parts: downsampling block, bottleneck block, and upsampling block. Each block consists of a combination of 2D convolutional layers, self-attention layers, and cross-attention layers.
During inference, a sample $\textbf{z}_{T}$ from a Gaussian distribution is gradually denoised using the trained denoising U-Net. Undergoing the denoising process from $t=T$ to $t=0$, the model generates $\textbf{z}_{0}$. This final latent representation is then passed through the decoder $\mathcal{D}$ to produce the output image.
Injecting human pose and camera viewpoint into diffusion model.
For conditions that should be spatially aligned with the generated output, we process them with a respective pose guider network and concatenate the processed conditions along the channel dimension. Specifically, the depth map $\textbf{d}\in\mathbb{R}^{H\times W\times 1}$, surface normals $\textbf{n}\in\mathbb{R}^{H\times W\times 3}$, and rendered human skeleton $\textbf{s}\in\mathbb{R}^{H\times W\times 3}$ are each processed by a respective pose guider network. This network reduces the spatial size of the condition to $1/8$ of the original size and embeds the pose information into $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times C}$ embeddings, aligning it with the size of the latent representation in the diffusion model. The last layer of the pose guider network is initialized to zeros to minimize the initial degradation during the fine-tuning stage of the pre-trained SD model. The processed conditions are then concatenated along the channel dimension and added to the projected noise before it is fed into the U-Net.
Injecting identity into diffusion model as a condition.
For human identity, unlike the condition that is spatially aligned with the output, it is not necessarily aligned with the output. In this regard, instead of adding the condition pixel-wise, we provide the identity information to the denoising U-Net with attention. Specifically, we design a reference U-Net that has the same architecture as the denoising U-Net, while the weights of the U-Net are initialized with a pre-trained Stable Diffusion model. To inject the identity of an image into the denoising U-Net, we first feed the image into the reference U-Net. Then, the identity information is shared with the denoising U-Net using self-attention for each block. In more detail, given the intermediate feature map from the denoising U-Net $f_{1}\in\mathbb{R}^{(h\times w)\times c}$ and from the reference U-Net $f_{2}\in\mathbb{R}^{(h\times w)\times c}$, they are concatenated along the spatial dimension, followed by the self-attention layer. Then, the first half of the output is used as the input for the following layers in the denoising U-Net. In this way, the two parallel branches can benefit from the extensive pre-trained knowledge of Stable Diffusion. Additionally, the identical architecture of the two branches facilitates training by sharing the same feature space. For the cross-attention part in Stable Diffusion where text embeddings from CLIP are used, we instead utilize image embeddings from the CLIP image encoder. This is possible because both text and image embeddings are trained to reside in the same embedding space.
Additional qualitative results
Figure 8: Additional qualitative results. Examples of generated images from the Pose-dIVE augmented datasets. The results demonstrate realistic rendering while preserving the identity of the reference images and aligning accurately with the target poses.
In the main paper, we present generated images of Pose-dIVE augmented datasets using MSMT17 and Market-1501, showcasing the effectiveness of our approach. In this section, we provide additional qualitative examples from the MSMT17, Market-1501, CUHK03 (D), and CUHK03 (L) datasets.
The generated images in the Pose-dIVE augmented datasets exhibit several key strengths. Firstly, they maintain a high level of realism, effectively capturing the nuances of natural human appearances. Secondly, successfully preserves the identity of individuals from the input reference images, ensuring that the augmented data remains faithful to the original person’s characteristics. This is particularly important for re-identification tasks where identity preservation is critical. Thirdly, the generated images accurately follow the provided target poses, demonstrating the robustness in handling diverse and complex human poses. These additional results further validate the effectiveness of our proposed augmentation method across different datasets.
Analysis on the number of generated images.
Figure 9: Impact of the number of generated images per PID. Experiments are conducted in the Pose-dIVE augmented CUHK03 (L) dataset. We use CLIP-ReID baseline.
Figure 9 illustrates the effect of varying the number of generated images on the performance of the Re-ID model when using our proposed augmentation strategy. To evaluate this, we gradually increased the number of generated images in the training dataset and trained the CLIP-reID model, carefully observing how performance changed with each increment. It’s important to emphasize that this augmentation is only applied to the training dataset, with the test dataset remaining unchanged throughout the experiments.
The analysis reveals a clear trend: there is an optimal number of generated images that maximizes the performance of the Re-ID model. This trend was consistently observed across all the datasets we analyzed, allowing us to fix the ratio of generated to original images regardless of the specific dataset.
Visualizing Conditioning in Generative Models
Figure 10: Example SMPL, skeleton, depth and normal maps from external dataset. Examples of generated images from the Pose-dIVE augmented datasets. The results demonstrate realistic rendering while preserving the identity of the reference images and aligning accurately with the target poses.
We present visualizations of conditions used in our generative model, specifically, the skeleton, depth, and normal maps, in Figure 10. With these conditions, we can control the human pose and viewpoint of the images generated by our diffusion models.